OpenStudy (anonymous):

Distribute. 5x = 21 - 3x

Parth (parthkohli):

If you mean FACTOR, then: \(\Large \color{MidnightBlue}{\Rightarrow 5x = 3(7 - x) }\)

OpenStudy (anonymous):

5x+3x=21 8x=21 x=21/8

Parth (parthkohli):

Or if you mean SOLVE: \(\Large \color{MidnightBlue}{\Rightarrow 8x = 21 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = {21\over 8} }\)

OpenStudy (anonymous):

5x=3(7-x)

OpenStudy (anonymous):

i got 5x = 21 - x on that one